OpenStudy (anonymous):

How many grams of a stock solution that is 87.0 percent H2SO4 by mass would be needed to make 200 grams of a 35.5 percent by mass solution? Show all of the work needed to solve this problem.

OpenStudy (photon336):

you weren't given any more information right?

OpenStudy (anonymous):

thats all that was given to me

OpenStudy (anonymous):

What do you know in general though?

OpenStudy (anonymous):

Are you adding water to it or something? What is the percent of the solution you're adding to it?

OpenStudy (photon336):

No concentration was given? usually the stock solution has a concentration

OpenStudy (photon336):

Ah i see

OpenStudy (anonymous):

My only guess is that you're adding water to it, or something with 0% solution. In that case \(s\) is grams of solution, and \(w\) is grams of water.\[ s+w=200\\ 0.87s + 0w = 0.355(s+w) \]

OpenStudy (anonymous):

My top equation is just the total grams equation. My bottom equations is the total \(H_2SO_4 \) grams equation.